OpenStudy (anonymous):

What is an equation in slope-intercept form for the line that passes through the points (1, –3) and (3, 1)? A. y=3x+1 B.y=x-3 C.y=2x+5 D. y =2x-5

OpenStudy (anonymous):

There will be 2 equations. Here's to get you started. y=mx+b Point (1,-3) -3=m(1)+b Point (3,1) 1=m(3)+b I think it's easiest to use the substitution method.

OpenStudy (anonymous):

Do you know how to do solve for it?

OpenStudy (anonymous):

yea can i give you what i think it is and you see if it right /

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

You're very close.

OpenStudy (anonymous):

not c

OpenStudy (anonymous):

yes! you can also check your answers by graphing them here's a link to a graphing calculator http://my.hrw.com/math06_07/nsmedia/tools/Graph_Calculator/graphCalc.html

OpenStudy (anonymous):

Thanks Alot , can i ask you more and you check my answers !

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Suppose y varies directly with x, and y = 24 when x = 8. What is the value of y when x = 10

OpenStudy (anonymous):

a.10 B.30 C.80 D240

OpenStudy (anonymous):

@helplesspenguin and i think it is 30

OpenStudy (anonymous):

you're right.good job

OpenStudy (anonymous):

any other questions?